A concave mirror has a radius of curvature of 1.35 m. A bouquet of roses is placed 1.24 m from the mirror on its reflecting side. At what distance from the mirror will the roses will be focused sharply on a screen? Answer in units of m.
Answer #1

1/p + 1/q = 1/f

f = R/2 = 1.35/2

1/1.24 + 1/q = 1/(1.35/2)

q=1.48 m